OpenStudy (anonymous):

solve the initial value dy/dt=(t^2)tan(y) y(0)=0

OpenStudy (lgbasallote):

this can be separated \[\implies \frac{dy}{\tan y} = t^2 dt\] does that help?

OpenStudy (anonymous):

Ive already done that so the integrals are ln/sin(y)/ = (t^3)/3+C right?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

y(0) = 0 so change t to 0 and y to 0

OpenStudy (lgbasallote):

then solve for x

OpenStudy (anonymous):

so when you plug in y(0) = 0 the left side doesnt work

OpenStudy (lgbasallote):

hmm interesting

OpenStudy (anonymous):

because you get ln of 0 which is undefined

OpenStudy (anonymous):

any ideas?

OpenStudy (lgbasallote):

maybe c just doesnt exist?

OpenStudy (anonymous):

okay do this \[\sin y=e ^{t ^{3}/3}+c\] now solve for c by putting y=0

OpenStudy (anonymous):

so c=1?

OpenStudy (anonymous):

sorry it will be \[e ^{c}\]

OpenStudy (lgbasallote):

\[\huge y = \sin^{-1} \left( e^{\frac{t^3} 3 + c_1} \right)\] nothing seems to fit

OpenStudy (anonymous):

@Igbasallote is my answer correct or will it be only c

OpenStudy (lgbasallote):

should be e^c

OpenStudy (lgbasallote):

but you can just call e^c as K

OpenStudy (lgbasallote):

then solve for K

OpenStudy (lgbasallote):

it's solveable

OpenStudy (anonymous):

dont you get e^c=0

OpenStudy (lgbasallote):

yes. e^c = 0

OpenStudy (lgbasallote):

so you can just put that...

OpenStudy (lgbasallote):

there's no expoennt you can raise e that will result to 0

OpenStudy (lgbasallote):

so expressing it as e^c maybe the best you can hope for

OpenStudy (anonymous):

true

OpenStudy (anonymous):

i can just say e^c is an arbitrary constant

OpenStudy (lgbasallote):

oh wait no

OpenStudy (lgbasallote):

it's really impossible lol

OpenStudy (lgbasallote):

the GS is \[\huge \sin y = e^{\frac{t^3}3} \times e^c\] so if e^c is 0...this becomes 0....

OpenStudy (anonymous):

thats what i thought

OpenStudy (lgbasallote):

might be discontinuous at y(0) = 0?

OpenStudy (anonymous):

true i didnt think of that

OpenStudy (anonymous):

\[y(0)=0\implies y'(0)=t^2\tan(y)=(0)(0)=0\]No change in \(y\) around the point \((0,0)\), so \(y\) remains constant. Check this using a slope field:So \(y=0\) is the solution.